Consider the following questions about buffers and determine if each statement is true or false.


true false  A buffer based on 1.0 M HF and 1.0 M NaF will be more acidic than a buffer based on 1.0 M acetic acid and 1.0 M sodium acetate. (Ka HF = 7.2 x 10-4, Ka acetic acid = 1.8 x 10-5)
true false  A 1.0 M solution of HF will be more acidic than a buffer comprised of 1.0 M HF and 1.0 M NaF
true false  A solution that contains equal concentrations of HOCl (a weak acid) and NaOCl will be a buffer
Answer #1

(1.) A buffer based on 1.0 M HF and 1.0 M NaF will be more acidic than a buffer based on 1.0 M acetic acid and 1.0 M sodium acetate. : True

(2.) A 1.0 M solution of HF will be more acidic than a buffer comprised of 1.0 M HF and 1.0 M NaF : True

(3.) A solution that contains equal concentrations of HOCl (a weak acid) and NaOCl will be a buffer : True

Explanation

(1.) Ka HF = 7.2 x 10-4

pKa = -log(Ka)

pKa = -log(7.2 x 10-4)

pKa = 3.14

According to Henderson - Hasselbalch equation,

pH = pKa + log([conjugate base] / [weak acid])

pH = pKa + log([NaF] / [HF])

pH = 3.14 + log(1.0 M / 1.0 M)

pH = 3.14 + log(1)

pH = 3.14 + 0

pH = 3.14

Ka acetic acid = 1.8 x 10-5

pKa = -log(Ka)

pKa = -log(1.8 x 10-5)

pKa = 4.74

According to Henderson - Hasselbalch equation,

pH = pKa + log([conjugate base] / [weak acid])

pH = pKa + log([acetate] / [acetic acid])

pH = 4.74 + log(1.0 M / 1.0 M)

pH = 4.74 + log(1)

pH = 4.74 + 0

pH = 4.74

Since pH for NaF / HF buffer is lower pH of acetate / acetic acid buffer, therefore, NaF / HF buffer is more acidic which means the given statement is true.